Please Help Me Figure This Out!!!!!!!!
Question- A baseball park has 40,000 seats of which 13,040 are box seats and 18,200 are reserved seats. What percent of the seats are box seats? What percent of the seats are reserved seats?
: Please Help Me Figure This Out!!!!!!!!
Question- A baseball park has 40,000 seats of which 13,040 are box seats and 18,200 are reserved seats. What percent of the seats are box seats? What percent of the seats are reserved seats?

Answer by jim_thompson5910(9166)


                         # of box seats
Percent of Box Seats = ----------------- * 100
                        total # of seats



                               # of reserved seats
Percent of Reserved Seats = ----------------------- * 100
                               total # of seats



Percent of Box Seats = (13040/40000)*100=0.326*100 = 32.6 %


Percent of Reserved Seats = (18200/40000)*100=0.455*100 = 45.5 %


Note: the percentages normally add to 100%, but in this case, there are more seats that are not considered.
Question 152848: Please Help Me Figure This Out!!!!!!!!
Question- A baseball park has 40,000 seats of which 13,040 are box seats and 18,200 are reserved seats. What percent of the seats are box seats? What percent of the seats are reserved seats?
: Please Help Me Figure This Out!!!!!!!!
Question- A baseball park has 40,000 seats of which 13,040 are box seats and 18,200 are reserved seats. What percent of the seats are box seats? What percent of the seats are reserved seats?

Answer by Fombitz(1740)

Per cent is the ratio out of 100.
Percentage of box seats to total seats is
x/100=13040/40000
x=(13040/40000)*100
x=32.6
Percentage of reserved seats to total seats is
x/100=18200/40000
x=(18200/40000)*100
x=45.5
32.6% of the total seats are box seats, 45.5% are reserved seats.